OpenStudy (anonymous):

write a function A polynomial with rational coefficients having roots 3, 3, and 3-i

OpenStudy (campbell_st):

well it will have a conjugate pair for 3 - i the polynomial will be P(x) = (x -3)(x -3)(x -(3 -i))(x +(3 - i)) just distribute for the answer

OpenStudy (cookiimonster627):

would it be( x-3^2)(x-(3-i))(x+(3-i)? @campbell_st

OpenStudy (campbell_st):

well complex solutions x = 3 - i subtract 3 from both sides (x - 3) = -i now square both sides \[(x - 3)^2 = (-i)^2\] and knowing i^2 = -1 you get \[x^2 - 6x + 9 = -1\] so the equadraict fractor is \[x^2 - 6x + 10 = 0\] and knowing the rational zeros you can say \[P(x) = (x - 3)(x -3)(x^2 - 6x + 10)\]

OpenStudy (cookiimonster627):

I got x^4-12x^3+55x^2-114x+90 would that be right or wrong? @campbell_st